OpenStudy (anonymous):

what is the derivative to ((3x^2)(e^2x-1))/((2x^2)(-3x) +1)). If you can explain how you got the answer it would be super helpful. Thanks

OpenStudy (wmj259):

First you have the use the quotient rule. At the same time you need to use the multiplication rule also.

OpenStudy (wmj259):

Product Rule I meant.

OpenStudy (freckles):

\[y=\frac{3x^2(e^{2x}-1)}{2x^2(-3x+1)}\]?

OpenStudy (wmj259):

So the quotient rule, how I remember it is Low-Di-High minus Hi-Di-Low all over Low*Low.

OpenStudy (freckles):

if you want to you could find the derivative using log implicit differentiation,

OpenStudy (freckles):

or actually you could cancel those x^2's out first

OpenStudy (freckles):

that would make the quotient rule prettier

OpenStudy (freckles):

But I'm not totally sure what I'm seeing exactly

OpenStudy (jdoe0001):

\(\bf \large { \cfrac{d}{dx}\left[\cfrac{(3x^2)(e^{2x-1})}{(2x^2)(-3x)}+1\right]? \ or \ \cfrac{d}{dx}\left[\cfrac{(3x^2)(e^{2x-1})}{(2x^2)(-3x+1)}\right]? }\)

OpenStudy (jdoe0001):

you could use the whiteboard, or post a picture btw

OpenStudy (anonymous):

\[\frac{ 3x ^{2}(e ^{2x-1})}{ (2x ^{2})(-3x+1)}\] Thanks

OpenStudy (anonymous):

Cancel by x^2 before doing the derivative

OpenStudy (anonymous):

\[ \frac{3 x^2 e^{2 x-1}}{2 x^2 (1-3 x)}=\frac{3 e^{2 x-1}}{2 (1-3 x)} \] Now use the quotient rule of the RHS

OpenStudy (anonymous):

Use this to check your answer \[ \frac {d}{dx} \frac{3 e^{2 x-1}}{2 (1-3 x)}=-\frac{3 e^{2 x-1} (6 x-5)}{2 (3 x-1)^2} \]

OpenStudy (anonymous):

Thank you!